# SPECIAL RELATIVITY

### v9

@@ 7.3 Acceleration of a Rigid Body and Special Relativity
Now let us look at some further consequences of relativity of simultaneity. The concept of rigid motion, as known from Newtonian mechanics, cannot be taken over unchanged to the special theory of relativity. This follows from the fact that in the space measurement of one observer there is mixing of space and time, as seen by other.
First, let us study the relativistic kinematics of rotation of the moving frame axes with respect to the inertial frame. Consider an emitter-detector setup in the lab inertial frame, Fig. 36. Suppose that the setup is accelerated from the rest with respect to the lab frame up to velocity $v$ along the $x$-axis. The main difference between the relativistic kinematics and the kinematics of Newton is that in the lab frame, the accelerated emitter-detector setup exhibits a shearing motion. This counterintuitive fact is not a paradox. In fact, this shearing contribution adds up to the Wigner rotation. The shearing (3-space geometric) effect is closely associated with the relativity of simultaneity.
@@
Above we demonstrated that aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, may be derived from the Lorentz transformations. In other words, aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, presents a kinematics effect of special theory of relativity, Fig. 37. It is shown that the relativity of simultaneity is responsible for aberrations in the first order of $v/c$. We discovered of the essential asymmetry between the inertial and accelerated frame, namely, an acceleration (with respect to the inertial frame) have an effect on the propagation of light in the accelerated frame. On accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter, Fig. 37. In contrast, the inertial detector at rest continues to receive light radiated by the inertial emitter at rest, Fig. 36 left.
Figure 36: Aberration of light in an inertial frame of reference. The aberration of light effect can be described within the standard special relativity taking advantage of the Wigner rotation theory. When one has a transversely moving emitter there is the deviation of the energy transport for radiated light. This effect is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves (see Fig. 1). As viewed from the Lorentz lab frame, the coordinate axes of the accelerated frame is rotated through the angle $v/c$ with respect to the coordinate axes of the lab frame. In the first order approximation in $v/c$, the accelerated emitter-detector setup exhibits a shearing motion (during acceleration). This shearing contribution adds up to the Wigner rotation. The wavefront orientation has no exact objective meaning. Because of this, only the wavefront orientation in the absolute time coordinatization for the moving source is presented here for sketch simplification.
@@
Figure 37: Aberration of light in an accelerated frame of reference. In order to predict result of the aberration measurement, the accelerated observer should use the Langevin metric. On accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter.
@@
Figure 38: Aberration of light in an inertial frame of reference as predicted by the conventional theory (see Fig. 1).
@@
Figure 39: Aberration of light in an accelerated frame of reference as predicted by the conventional theory. An acceleration does not spoil the motional symmetry between the accelerated reference frame and inertial reference frame.
@@
Common textbook presentations of the special relativity use the (3+1) approach which deals with hidden assumption that distant clocks are synchronized according to the absolute simultaneity convention. In fact, according to textbooks, when a reference frame at rest is put into motion, then all points in the 3D reference grid have a start to move at same time. Consequently, rigidity in the sense introduced by textbooks cannot be considered as a relativistic kinematics property. The standard presentation of the aberration of light effect is based on the hidden assumption that $(x_{x},y_{n},z_{n})$ axes of the accelerated frame and $(x,y,z)$ axes of the inertial frame are parallel. In other words, it is incorrectly assumed that accelerated and inertial observers have common 3-space, Fig. 38,39. According to the conventional theory, an acceleration does not spoil the motional symmetry between the accelerated reference frame and inertial reference frame. Our result (Fig. 36,37 ) is at odds with the prediction from textbooks. The commonly accepted derivation of the aberration of light effect does not account for the Wigner rotation, which in our case closely associated with the relativity of simultaneity.
@@ 7.4 Discussion
Let us return to observations of an inertial observer. Special theory of relativity says that the orthogonal emitter-detector setup exhibits a shearing motion during acceleration, Fig. 36. This is a purely kinematic effect, involving no forces. How can that be? That’s not an easy question, but here is one way of thinking about this.
@@
It is clear that our ordinary intuition are quite wrong when we have deal with two space-like separated events. We see that the accepted in textbooks assumption of rigidity of accelerated frame is based on the belief that simultaneous acceleration of the reference space grid has direct physical meaning. This is misconception. In fact, when we have two distant events, we have deal with the conventionality of distant simultaneity within the time interval $L/c$, where $L$ is the space separation of these two events. The position along the $x$-axis of the accelerated emitter with respect to the accelerated detector has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained . There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $Lv/c$ (due to the uncertainty in the moments of the acceleration).
@@
There are several points to be made about the above results. Above we discussed that the aberration of light effect in accelerated frames can be described in an inertial frame taking advantage of the Wigner rotation theory. To be complete we must add an analysis of the vertical displacement of the detector. We have considered only the case in which motion is non-relativistic in speed. In other words, we use a first order relativistic approximation. It is easy to see that for the vertical displacement we have $\Delta x\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$. So special relativity says that there is no difference between the (kinematic) shearing motion and the (Wigner) rotation motion (during acceleration) as $v^{2}/c^{2}\to 0$.
7.5 Aberration of Relativistic Particles
We want now to discuss the phenomenon of the aberration of relativistic particles. We begin by recalling a very interesting (methodological) analogy between the aberration of light effect and the deflection of light in general relativity. Consider a ”point” mass $m$ moving with a velocity $V$ in the gravitation field of a point mass $M$. Both the motion of particles and the path of light in the gravitation field are described by the same equation of motion. The deflection angle for a particle is $\theta_{d}=(1+V^{2}/c^{2})\theta_{c}$, where $\theta_{c}$ is the particle deflection angle according to classical approach. The factor $(1+V^{2}/c^{2})$ is new from relativistic (Schwarzschild) metric, but reduced to familiar Newtonian limit as $V^{2}/c^{2}\to 0$. For light $V/c=1$ and there is an extra factor 2. On the basis of classical theory one gets a deflection of light that is only one-half as big as predicted by general relativity.

### v10

@@ 7.3 Acceleration of a Rigid Body and Special Relativity
Now let us look at some further consequences of relativity of simultaneity. The concept of rigid motion, as known from Newtonian mechanics, cannot be taken over unchanged to the special theory of relativity. First, let us study the relativistic kinematics of rotation of the moving frame axes with respect to the inertial frame axes. Consider two emitter-detector setups in the initial inertial frame. Suppose one setup is accelerated from the rest with respect to the inertial frame up to velocity $v$ along the $x$-axis, Fig. 32. The explanation of the effect of aberration is based on the use of absolute time coordinatization  [^36] Another explanation consists of using a clock re-synchronization procedure in the initial inertial frame. The time under Einstein’s synchronization for the moving source is readily obtained by introducing the offset factor $xv/c^{2}$ in the first-order approximation. This time shift has the effect of rotation of the radiation wavefront on the angle $v/c$ for both sources. In this synchronization, we describe the radiation of the moving source using Maxwell’s equations. It is interesting to discuss the electrodynamics of the source at rest. The new crossed term that has to be put into the wave equation due to time transformation leads to the prediction of the group velocity $d\omega/dk_{x}=-v$ (see footnote in Section 3.5). On the other hand, there is the angle $v/c$ between the wave number vector $\vec{k}$ and the $z$-axis. Then, accounting for the initial condition, the radiated light beam of the source at rest moves along the $x$-axis with group velocity $d\omega/dk_{x}=0$ as it must be.. The main difference between the relativistic kinematics and the kinematics of Newton is that in the inertial frame, the accelerated emitter-detector setup exhibits a shearing motion. This counterintuitive fact is not a paradox. The shearing (3-space geometric) effect is closely associated with the relativity of simultaneity.
@@
Above we demonstrated that the aberration of light effect in the accelerated frame, as viewed from the inertial frame, may be derived from the Lorentz transformations. In other words, aberration of light effect in the accelerated frame, as viewed from the initial inertial frame, presents a kinematics effect of the special theory of relativity, Fig. 33. It is shown that the relativity of simultaneity is responsible for aberrations in the first order of $v/c$. We discovered the essential asymmetry between the inertial and accelerated frame, namely, an acceleration (with respect to the inertial frame) has an effect on the propagation of light in the accelerated frame. In an accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter, Fig. 33. In contrast, the detector at rest in the inertial frame continues to receive light, Fig. 32 left.
Figure 32: Aberration of light in an inertial frame of reference. The wavefront orientation has no exact objective meaning. Because of this, only the wavefront orientation in the absolute time coordinatization for the moving source is presented here for sketch simplification.
@@
Figure 33: Aberration of light in an accelerated frame of reference. In an accelerated system, the accelerated detector does not receive light radiated by the accelerated emitter.
@@
Figure 34: Aberration of light in an inertial frame of reference as predicted by the conventional theory.
@@
Figure 35: Aberration of light in an accelerated frame of reference as predicted by the conventional theory. Acceleration does not spoil the motional symmetry between the accelerated reference frame and the inertial reference frame.
@@
Common textbook presentations of the special relativity use the (3+1) approach which deals with the hidden assumption that distant clocks are synchronized according to the absolute simultaneity convention. According to textbooks, when a reference frame at rest is put into motion, then all points in the 3D reference grid start to move at the same time. The standard presentation of the aberration of light effect is based on the hidden assumption that $(x_{x},y_{n},z_{n})$ axes of the accelerated frame and $(x,y,z)$ axes of the inertial frame are parallel. In other words, it is incorrectly assumed that accelerated and inertial observers have common 3-space, Fig. 34 - Fig.35. According to the conventional theory, acceleration does not spoil the motional symmetry between the accelerated reference frame and the inertial reference frame. Our result (Fig. 32 - Fig.33 ) is at odds with the prediction from textbooks. The commonly accepted derivation of the aberration of light effect does not account for the Wigner rotation, which in our case is closely associated with the relativity of simultaneity. In the next chapter we shall consider the Wigner rotation theory in great detail.
@@ 7.4 Discussion
Let us return to observations of an inertial observer. The special theory of relativity says that the orthogonal emitter-detector setup exhibits a shearing motion during acceleration, Fig. 32. This is a purely kinematic effect, involving no forces. How can that be? That’s not an easy question, but here is one way of thinking about this.
@@
It is clear that our ordinary intuition is quite wrong when we have to deal with two space-like separated events. We see that the generally accepted assumption of rigidity of the accelerated frame is based on the belief that simultaneous acceleration of the reference space grid has direct physical meaning. This is a misconception. When we have two distant events, we have to deal with the conventionality of distant simultaneity within the time interval $L/c$, where $L$ is the space separation of these two events. The position along the $x$-axis of the accelerated emitter with respect to the accelerated detector has no exact objective meaning since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this position could be ascertained. There is an uncertainty (blurring) of the relative position in the $x$-direction of amount $Lv/c$ (due to the uncertainty in the moments of the acceleration)  [^37] Our next example has to do with a light clock. A clock which we call a ”light clock” is a rod with a mirror at each end. According to textbooks, when the inertial observer looks at the orthogonal light clock going by, he sees that light, in going from mirror to mirror is taking a zigzag path. This standard explanation is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving observer and $(x,y,z)$ axes of the inertial observer are parallel. This is a misconception. Let us consider the phase (and initialization) of light clocks. The word ”phase” means the zigzag shape of the path followed by the light beam. The shape of the zigzag pattern cannot be measured in the lab frame. The position of the accelerated bottom mirror with respect to the accelerated top mirror has no exact objective meaning. There is uncertainty of amount $lv/c$, where $l$ is the rod length..
@@
There are several points to be made about the above results. We already know that the aberration of light effect in accelerated frames can be described in an inertial frame taking advantage of the Wigner rotation theory. One of the important conclusions of the discussion presented above is that inertial and accelerated observers have no common line of motion. To be complete we must add an analysis of the angular displacement of the $x_{n}$-axis. We have considered only the case in which motion is non-relativistic in speed. In other words, we use a first-order relativistic approximation. It is easy to see that for the displacement of the detector along the $x$-axis in the $z$ direction we have $\Delta z\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$. So special relativity says that there is no difference between the (kinematic) shearing motion (during acceleration) and the (Wigner) rotation motion as $v^{2}/c^{2}\to 0$.
8 Kinematics of the Wigner Rotation
The subject of this section is the Wigner rotation  [^38] As known, a composition of noncollinear Lorentz boosts does not result in a different boost but in a Lorentz transformation involving a boost and a spatial rotation, the latter being known as Wigner rotation [39, 40]. Wigner rotation is sometimes called Thomas rotation (see e.g. [13, 41]).. We have already discussed the Wigner rotation in the previous section. Our earlier discussion is really about the aberration of light. The Wigner rotation is also related to the aberration of particles. The Wigner rotation is a relativistic kinematic effect, which consists in that the coordinate axes of a reference frame, moves along a curvilinear trajectory rotating about the axes of a Lorentz inertial frame. To deduce this expression is the main purpose of this chapter. We have tried to keep the mathematical complexity of the discussion to a minimum.
8.1 The Commutativity of Collinear Lorentz Boosts
Let us now consider a relativistic particle, accelerating in the initial inertial frame, and let us analyze its evolution within Lorentz coordinate systems. The permanent rest frame of the particle is not inertial. To get around that difficulty one introduces an infinite sequence of comoving frames. At each instant, the rest frame is a Lorentz frame centered on the particle and moving with it. As the particle velocity changes to its new value at an infinitesimally later instant, a new Lorentz frame centered on the particle and moving with it at the new velocity is used.
Let us denote the three inertial frames by $K,K^{\prime},K"$. The initial inertial frame is $K$, $K^{\prime}$ is the particle’s rest frame with velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and $K"$ is the particle’s rest frame at the next instant of proper time $\tau+d\tau$, which moves relative to $K^{\prime}$ with infinitesimal velocity $d\vec{v^{\prime}}$. All inertial reference frames are assumed to be Lorentz reference frames. To have this, we impose that $K^{\prime}$ is connected to $K$ by the Lorentz boost $L(\vec{v})$, with $\vec{v}$, which transforms a given four vector event $X$ in space-time into $X^{\prime}=L(\vec{v})X$. The relation $X"=L(d\vec{v^{\prime}})L(\vec{v})X$ presents a step-by-step change from $K$ to $K^{\prime}$ and then to $K"$.
There is another composition of reference-frame transformations that describes the same particle evolution in the Minkowski space-time. Let $R^{\prime}$ be an inertial frame with velocity $d\vec{v}$ relative to the initial inertial frame $R$. We impose that $R^{\prime}$ is connected to $R$ by the Lorentz boost $L(d\vec{v})$. The Lorentz rest frame $R"$ is supposed to move relative to the Lorentz frame $R^{\prime}$ with velocity $\vec{v}$. The relation $X"=L(\vec{v})L(d\vec{v})X$ presents a step-by-step change from $R$ to $R^{\prime}$ and then to the rest frame $R"$.
Let us examine the transformation of the three-velocity in the theory of relativity. For a rectilinear motion along the $z$ axis, it is performed in accordance with the following equation: $v_{z}(\tau+d\tau)=(dv_{z}+v_{z})/(1+v_{z}dv_{z}/c^{2})$. As it happens with the composition of Galilean boosts, collinear Lorentz boosts commute: $L(dv_{z})L(v_{z})=L(v_{z})L(dv_{z})$. This means that the resultant of successive collinear Lorentz boosts is independent of which transformation applies first.
In contrast with the case of Lorentz boosts in collinear directions, Lorentz boosts in different directions do not commute. While the successive application of two Galilean boosts is a Galilean boost and the successive application of two rotations is a rotation, the successive application of two non-collinear Lorentz boosts is not a Lorentz boost.
Let us compare the succession $K\to K^{\prime}\to K"$ with the succession $R\to R^{\prime}\to R"$ in the case when the particle’s acceleration in the frame $K^{\prime}$ is perpendicular to the line of flight of the initial inertial frame. The frame $K"$ is supposed to move relative to $K^{\prime}$ with velocity $d\vec{v^{\prime}}_{x}$. Now, since we can write the result in terms of succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ as well as in terms of succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$, there is a need to clarify a number of questions associated with these compositions of Lorentz frames. We can easily understand that the operational interpretation of the succession $L(\vec{v}_{z})L(d\vec{v}_{x})$ is particularly simple, involving physical operation used in the measurement of the particle’s velocity increment $d\vec{v}_{x}$ in the initial inertial frame. We should be able to understand the operational interpretation of the succession $L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})$. The inertial observer sees the time dilation in the Lorentz frame which moves with respect to the initial inertial frame with velocity $\vec{v}_{z}$: $dt/\gamma=d\tau$. Because of time dilation in the moving frame, the velocity increment in the initial inertial frame $dv_{x}$ corresponds to a velocity $\gamma dv_{x}$ and $-\gamma dv_{x}$ in the frames $K^{\prime}$ and $K"$ respectively. The resulting boost composition can be represented as $X"=L(d\vec{v^{\prime}}_{x})L(\vec{v}_{z})X=L(\vec{v}_{z})L(d\vec{v}_{x})X$. In other words, Lorentz boosts in different directions do not commute: $L(\vec{v}_{z})L(d\vec{v}_{x})\neq L(d\vec{v}_{x})L(\vec{v}_{z})$.
8.3 Wigner Rotation
[…]
There are several points to be made about the Wigner rotation theory. We wish to remark that the expression for the Wigner rotation angle in the proper frame (as viewed from the Lorentz $K$ frame), $d\vec{\Phi}_{K"}=(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$, can be presented in the form $d\vec{\Phi}_{K"}=(1-1/\gamma)\vec{v}_{K"}\times d\vec{v}_{K"}/v_{K"}^{2}$. Thus, the Wigner rotation angle in the proper frame is expressed in terms of the $K$ frame velocity $\vec{v}_{K"}=-\vec{v}$ and its increment $d\vec{v}_{K"}=-\gamma d\vec{v}$ in the proper frame.
Transformed back to the lab Lorentz frame this becomes $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$. This is the Wigner rotation angle of the proper frame axes in the initial inertial (lab) frame $K$  [^39] The correct expression for the Thomas (Wigner) rotation was first obtained by V. Ritus [42]. In deriving expressions for the Thomas (Wigner) rotation, the majority of authors (see e.g. [41]) was supposedly guided by the incorrect expression for Thomas (Wigner) rotation from Moeller’s monograph [13]. The expression obtained by Moeller is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$ (and subsequently expression for Thomas precession $\Omega_{\mathrm{T}}=(1-\gamma)\omega_{0}$). It should be noted that, in his monograph, Moeller stated several times that this expression is valid in the lab Lorentz frame. Clearly, this expression and correct result differ both in sign and magnitude. An analysis of why Moeller obtained an incorrect expression for the Thomas (Wigner) rotation in the lab frame is the focus of Ritus paper [43]. As shown in [43], the Moeller’s mistake is not computational, but conceptual in nature. In review [44] it is shown that the correct result was obtained in the works of several authors, which were published more than a half century ago but remained unnoticed against the background of numerous incorrect works.. We note that owing to the relativistic effect of time dilation in the reference frame that moves to the lab frame, the Wigner rotation angle in the proper frame is always $\gamma$ time higher than in the lab frame  [^40] We note that in 1986, M. Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [45]. It is noteworthy that [45] is one of the few papers that explicitly states that the angle of the Thomas (Wigner) rotation in the comoving reference frame is $\gamma$ times higher than in the lab frame..
We derived the exact relation $d\vec{\Phi}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ using only rudimentary knowledge of special relativity. In textbooks on the theory of relativity, the spatial rotation associated with the composition of two Lorentz boosts in non-parallel directions is often introduced using the algebraic approach. This is one of the reasons why authors of textbooks obtained an incorrect expression for the Wigner rotation. They describe the rotation of a moving object without the geometrical meaning of such rotation and encounter serious difficulties in the interpretation of the applied calculations and of the results.
Figure 36: The interpretation of the Wigner rotation about the proper frame axes.
8.4 How to Measure a Wigner rotation?
The Wigner rotation is a relativistic kinematic effect, which consists in that the coordinate axes of a reference frame, moving along a curvilinear trajectory rotating about the axes of a Lorentz lab frame. In the Lorentz lab frame, for infinitely small transformations (due to acceleration) we obtained the formula
where $d\vec{v}$ is the vector of small velocity change due to acceleration, $\Phi$ is the Wigner rotation angle of the spatial coordinate axes of the proper system relative to Lorentz lab frame, and $\theta$ is the orbital angle. But how to measure this orientation? A moving coordinate system changes its position in time. The question arises whether it is possible to give an experimental interpretation of the rotation of a moving coordinate system. We illustrate the problem of how to represent the orientation of the moving coordinate system with a simple example.
The execution of successive transformations from $R$ to $R^{\prime}$ at velocity $d\vec{v_{x}}$ and from $R^{\prime}$ to $R"$ at velocity $\vec{v_{z}}$ equivalent to the composition of boost $L(\vec{v}_{z}+d\vec{v}_{x})$ and rotation. The Wigner rotation Eq.(20) is performed additionally to the Lorentz boost at velocity $\vec{v_{z}}+d\vec{v_{x}}$. The interpretation of this rotation about the initial inertial frame of reference is closely associated with the length contraction  [^41] Length of moving object is determined not only by the properties of space-time but also by our synchronization convention. Therefore, similar to the radiation phase front orientation, it has no exact objective meaning. Such quantities (under the standard clock synchronization) usually are called ”relativistic kinematics effects”. The Wigner rotation is also a coordinate effect with no exact objective meaning..
Figure 37: The interpretation of the Wigner rotation about the lab frame.
Suppose that the lab observer after the Lorentz boost at velocity $v_{z}$ rotates the coordinate system on the angle $\vec{\delta\theta}=\vec{v_{z}}\times d\vec{v_{x}}/v_{z}^{2}$ and now $x_{r}$ locates orthogonally to the vector $\vec{v_{z}}+d\vec{v_{x}}$. Similarly, the axis $z_{r}$ is parallel to the vector $\vec{v_{z}}+d\vec{v_{x}}$. Consider a rod directed along $x$-axis in the comoving frame. The motion takes place in the plane $(x,z)$ and the rod is located perpendicularly to the velocity $\vec{v_{z}}$. After the rotation of the lab frame axes, the projection of the rod on the $z_{r}$ axes will be simply $l\delta\theta$, where $l$ is the rod length in the $R$ frame and also in the lab frame after the first boost along the $z$-axis. After the second Lorentz boost at velocity $d\vec{v_{x}}$ this projection will be contracted down to $l\delta\theta/\gamma$ (Fig. 37). Let the observer in the lab frame fix the position of the axes of the comoving frame. In ultrarelativistic limit $\gamma\to\infty$ these axes will be parallel to the rotated lab frame axes $(x_{r},z_{r})$. The projection of the rod on the $z_{r}$ axis will be zero. In the case of an arbitrary velocity, the axes of the comoving frame are not parallel to the rotated lab frame axes $(x_{r},z_{r})$. According to contracted projection, the angle will be $-\vec{\delta\theta}/\gamma$. And one can verify directly that the axes of the comoving frame are rotated with respect to the initial lab frame axes $(x,z)$ by the angle equal to $\delta\theta-\delta\theta/\gamma$, which is just the Wigner rotation angle in accordance with equation Eq.(20). Here we only wished to show how naturally Lorentz transformations lead to the Wigner rotation phenomenon.
8.5 Discussion
A large number of incorrect expressions for the Wigner (Thomas) rotation can be found in the literature. Let us discuss the reasons that underlie the derivation of the incorrect expressions. The Wigner rotation also relates the angular rotation velocity (Thomas precession) of the spin of elementary particles following a curvilinear orbit. In 1959, Bargman, Michel, and Telegdi (BMT) proposed a consistent relativistic theory for the dynamics of the spin as observed in the lab frame, which was successfully tested in experiments [46]. It is commonly believed that the BMT equation contains the standard (and incorrect) result for the Wigner rotation in the Lorentz lab frame. There are many physicists who have already received knowledge about the Wigner (Thomas) rotation from well-known textbooks and who would say, ”The extremely precise measurements of the magnetic-moment anomaly of the electron made on highly relativistic electrons are based on the BMT equation, of which the Wigner rotation is an integral part, and can be taken as experimental confirmations of the standard expression for the Wigner rotation”. This misconception about the experimental test of the incorrect expression for the Wigner rotation in the lab frame is common and pernicious. We should remark that the results in the Bargmann-Michel-Telegdi paper were obtained by the method of semi-classical approximation of the Dirac equation. The Wigner rotation was not considered at all, because the Dirac equation allows obtaining the solution for the total particle’s spin motion without explicitly splitting it into the Larmor and Wigner parts.
In should be noted that the situation relating to the use of the Wigner (Thomas) rotation theory in the relativistic spin dynamics is complex. The starting point of Bargman, Michel, and Telegdi was the particle rest frame and the equation of motion for particle angular momentum, which they generalized to the Lorentz lab frame and then transformed back to the rest frame. This brings up an interesting question: Why it is convenient to transform this equation to the rest frame as of that instant? The approach in which we deal with the proper spin is much preferred in the experimental practice due to the clear physical meaning of the (3D) spin vector $\vec{s}$. Unlike the (3D) momentum $\vec{p}$ , which has definite components in each reference frame, angular momentum is defined only in one particular reference frame. It does not transform. Any statement about it refers to the rest frame as of that instant. If we say that in the lab frame, the spin of a particle makes the angle $\phi$ with its velocity, we mean that in the particle’s rest frame, the spin makes this angle with the line motion of the lab frame. The analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard (Einstein) special relativity taking advantage of well-known relativistic kinematic effects. In the case of spin dynamics, one has no occasion to make measurements with non-inertial devices.
The conventional method used to explain the spin dynamics in the lab frame is very unusual. First, we already pointed out that the laws of physics in the lab frame are able to account for all physical phenomena, including the observations made by non-inertial observers. In particular, it is possible to determine the result of spin orientation measurements with respect to the lab frame axes in the proper frame. In this case, when viewed from the lab frame, the interpretation of the spin rotation experiments in the lab frame is simple and may be derived from the prediction of the proper observer as concerns the measurement made by the lab observer. Therefore, the spin rotation measurement by a polarimeter in the lab frame is interpreted from the viewpoint of the proper observer as viewing of the lab observer. This complexity is one of the reasons why authors of textbooks obtained an incorrect expression for the Wigner rotation. We will discuss this subject further in Chapter 17.
It is also important to elucidate the reason why did the error in the Wigner rotation theory remained so long undetected. Wigner (Thomas) rotation is generally presented only in the context of spin kinematics as a peculiar effect of special relativity. This is supposedly the reason why until recently no researches have noticed so serious a discrepancy between the results of different works.
We emphasize that the Wigner rotation is a very basic relativistic phenomenon. It is as basic as the time dilation and length contraction. The theory of special relativity shows that the aberration effect in the rotating earth-based frame, as viewed from the inertial sun-based frame (in the Lorentz coordinatization), presents a kinematic effect of the special theory of relativity. In this case, when viewed from the sun-based frame, the aberration image shift in the earth-based frame is regulated by the Wigner rotation. According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces. In contrast, the accepted in previous literature incorrect assumption that an earth-based observer and an inertial (e.g. sun-based) observer have a common 3-space has always been considered obvious.
9 Aberration of Relativistic Particles